What is the concentration of KOH if the solution has a pH of 11.56?
1
Expert's answer
2020-04-06T11:52:12-0400

pOH=14-pH

pOH=14-11.56=2.44

pOH=-lg(OH-)

2.44=-lg(OH-)

lg(OH-)=-2.44

(OH-)=10^(-2.44)

C(KOH)=0.00363